OpenStudy (anonymous):

Find the sum and choose the correct answer. 5n3 + 4n2 - 9 and -2n3 + 7n - 5

OpenStudy (mertsj):

Add the like terms 5n^3-2n^3=3n^3 4n62=4n^2 7n=7n -9+-5--14 So: 3n^3+4n^2+7n-14